OpenStudy (anonymous):

I want to check myself on this one: a/2 + 1/4 divided by a/3 - 1/2 you need same denominator so you multiply a/2 by 2 which gives you 2a/4 + 1/4 = 3a/4. Next part a/3 - 1/2 so you multiply a/3 by 2 to get 2a/6 then 1/2 you multiply by 3 to get 3/6 so 2a/6 - 3/6 = 2a-3/6. Now 3a/4 divided by 2a-3/6 then you use the reciprocal and multiply 3a/4 * 6/3 to get a/-2 (after you reduce and cancel out like terms)?

OpenStudy (asnaseer):

your 1st step has an error:\[\frac{a}{2}+\frac{1}{4}=\frac{2a}{4}+\frac{1}{4}\ne\frac{3a}{4}\]

OpenStudy (asnaseer):

do you understand why?

OpenStudy (anonymous):

Not sure .. now I'm confused LOL

OpenStudy (asnaseer):

look at the numerators of that fraction: 2a + 1 = ?

OpenStudy (anonymous):

it would just be 2a + 1 / 4

OpenStudy (anonymous):

because 2a and 1 are separate

OpenStudy (asnaseer):

yes - that is right

OpenStudy (anonymous):

So then it would be 2a + 1/4 divided by 2a - 3 / 6

OpenStudy (asnaseer):

your 2nd step was correct:\[\frac{a}{3}-\frac{1}{2}=\frac{2a-3}{6}\]

OpenStudy (asnaseer):

your 3rd step has an error

OpenStudy (anonymous):

then 2a + 1 / 4 * 6 / 2a - 3

OpenStudy (asnaseer):

what you want is:\[\frac{2a+1}{4}\div\frac{2a-3}{6}=\frac{2a+1}{4}\times\frac{6}{2a-3}\]yes - you have it right now :)

OpenStudy (anonymous):

yea I saw my error from the first part and realized I made the same mistake and then corrected it

OpenStudy (asnaseer):

ok - so you know what to do now and understand the process?

OpenStudy (anonymous):

Now I am unsure of the rest ...

OpenStudy (asnaseer):

:) np - just let me know what part is troubling you...

OpenStudy (anonymous):

you cannot do anything with the 2a + 1 or the 2a - 3 because the 2a's are attached to something, so the only thing you can mess with are the 4 and 6, correct?

OpenStudy (asnaseer):

yes

OpenStudy (anonymous):

(2a + 1)*3 / 2*(2a - 3) = 6a + 3 / 4a - 6 ?

OpenStudy (asnaseer):

that is correct. although I would leave it in the form:\[\frac{3(2a+1)}{2(2a-3)}\]

OpenStudy (anonymous):

haha yea i just noticed it was already reduced and doing what I did, UN-reduced it LOL

OpenStudy (asnaseer):

gr8! the good thing is that you recognised your mistakes - which is vital if you want to improve! :D

OpenStudy (anonymous):

I second guess myself all the time and change things when they don't need changed or over-think it!

OpenStudy (asnaseer):

he he - don't worry too much about it - the main thing here is that you have a good "hunger" for knowledge which means you will go far my friend! :D

OpenStudy (anonymous):

thanks for your help! (I haven't "math-ed" in about 10 years gosh, that probably gives away my age) I'm a college student again and I'm taking Technical Algebra ... I actually like math now! =)

OpenStudy (asnaseer):

I left college many many moons ago - oh no! - now I've given away my age as well :) Like you, I still have a thirst for knowledge. :D

OpenStudy (anonymous):

You're as "young" as you feel! take care and have a blessed day =) =)

OpenStudy (asnaseer):

thx - you too my friend! :)